OpenStudy (anonymous):

Finding y using the quadratic equation, y+16=225/y

OpenStudy (lieutenantgeneral):

y = 209 209 + 16 = 225

OpenStudy (anonymous):

No, no.. I have to arrange it like to this \[(y+16)/1-(225/y)=0\] and then use \[x=(-b±√(b^2-4ac))/2a\] , to find the solution(s) for y

OpenStudy (lieutenantgeneral):

OK.

OpenStudy (anonymous):

... yeah, how do i do that? how do I simplify it to insert it in that equation ?

OpenStudy (anonymous):

\[ y+16=\frac {225}y \\ y^2 + 16 y =225\\ y^2 + 16 y - 225 =0 \] Use now the quadratics formula

OpenStudy (campbell_st):

why not multiply every term by y and you get \[y^2 + 16y = 225\] now you have a quadratic that you can solve or \[y^2 + 16x -225 = 0\]

OpenStudy (anonymous):

Actually this factors to \[ (-9 + y) (25 + y) \]

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

y=9 y=-25

OpenStudy (anonymous):

oh... that seemed easy. lol So i just had to multiply y on both sides? thanks both @eliassaab and @campbell_st